OpenStudy (jotopia34):

What is the shortcut to finding the probability of having a certain outcome in biology genetics. It is supposedly solvable using a Prunett square which is a square showing all the possibilities but when you get to 16 total possibilities, I can't possibly draw a 16 answer box for every question. Others did this with 2 boxes of 4 each, but I can't figure out why. Help! The concrete example question is: Mother is YYRR and Father is yyrr. What are the chances of an offspring having a YRyr combination.

OpenStudy (anonymous):

Nice question.. a bit tricky too

OpenStudy (jotopia34):

They say that this is the same as rolling dice. Again, not seeing how..... :(

OpenStudy (jotopia34):

a concrete example is: Mother is YYRR and Father is yyrr. What are the chances of an offspring having a YRyr combination.

OpenStudy (anonymous):

so there possible value are Y ,y R,r ?

OpenStudy (anonymous):

first one you have 4 choice then 3 , so on 4*3*2*1=24

OpenStudy (jotopia34):

I understand how to find the total possible combinations but not the chance of getting a specific one. Also in YYRR and yyrr, the possibilities can be Yyrr, or YYrr, etc etc, but what are the chances of getting a an offspring that would have any order of capital Y and small r.... More specifically , a square would show all combos but i only filled in some below. I know there is a shorter way than writing this out and literally counting the number of times a specific one would show up for every question. I just can't figure out how. :( I hope I'm being remotely clear in the way I'm wording this.... YY RR YR RY yy YYyy RRyy YRyy RYyy rr yr r y

OpenStudy (anonymous):

if they have uniform chance, it would be 1/24

OpenStudy (jotopia34):

how did you get that number?

OpenStudy (anonymous):

see above

OpenStudy (jotopia34):

I'm sorry, I do see that, thank you but what I mean is, I thought there were only 16 possibilities?

OpenStudy (anonymous):

how did you see 16 possiblity

OpenStudy (jotopia34):

from the possibility square I drew, the prof drew that in class

OpenStudy (jotopia34):

The simpler version is mom has 2 genes for B, blue eyes (BB) and dad has 2 genes for brown (b) eyes. What are the chances of having an offspring with the Bb combination. mom B B Dad b Bb Bb b Bb Bb The chance is 100%, BUT, if mom has Bb and dad has bb, what are the chances of having bB offspring mom B b dad b Bb bb b Bb bb The this probability is 50% In the other question, the number of genes is doubled. Does this help?

OpenStudy (anonymous):

if mom is Bb and dad is bb we could have only two possibility first letter could either be B or b =2 second letter could just be b =1 2*1=2 2 choices 1/possible choices= probability 1/2=50% so 50% previous example BB and bb first letter could be B =1 seond letter could be b=1 1*1=1 so 100%

OpenStudy (anonymous):

clear?